OpenStudy (anonymous):

Solve y over negative 4 + 5 = 15. A.80 B.-80 C.40 D.-40

OpenStudy (ranga):

\[\Large \frac{ y }{ -4 } + 5 = 15 ?\]

OpenStudy (anonymous):

yes that's the question

OpenStudy (ranga):

To solve for y, we need to isolate y. Subtract 5 from both sides. Multiply both sides by -4 to find y.

OpenStudy (anonymous):

how do you isolate the y ?

OpenStudy (ranga):

I gave you the steps in line 2 and line 3 of my previous reply.

OpenStudy (anonymous):

ohok

OpenStudy (anonymous):

69875x45879

OpenStudy (anonymous):

please help mea

OpenStudy (anonymous):

what does this have to do with my question

OpenStudy (anonymous):

I really suck at algebra can you please explain it a little more @ranga

OpenStudy (ranga):

\[\frac{ y }{ -4 } + 5 = 15\]Subtract 5 from both sides.\[\frac{ y }{ -4 } = 10\]Multiply both sides by -4:\[y = 10 \times (-4) = -40\]

OpenStudy (anonymous):

I it just multiplying @germya

OpenStudy (anonymous):

is it B -80?

OpenStudy (anonymous):

please help @ranga im missing only 2 questions from my assignment

OpenStudy (anonymous):

well my bad do u still want help or nah

OpenStudy (anonymous):

HOW OLD IS YOU ANYWAY AND NO MY TUTOR JUST CAME OVA SO BYE!

OpenStudy (anonymous):

don't worry about my age and cool

OpenStudy (ranga):

The problem has been fully solved in my previous reply. How did you choose the wrong answer B?

OpenStudy (anonymous):

I honestly am very very bad at algebra like I barely understand it @ranga also can you help me with the last question ?

OpenStudy (anonymous):

@ranga